OpenStudy (anonymous):

If f(x)=√(x^3)-1, find the solution to the equation f(x)=f^-1(x).

Parth (parthkohli):

Let's say f(x) = y \(\Large \color{Black}{\Rightarrow y= \sqrt{x - 3} - 1 }\) \(\Large \color{Black}{\Rightarrow y +1 = \sqrt{x - 3} }\) \(\Large \color{Black}{\Rightarrow y^2 + 2y + 1 = x - 3 }\) \(\Large \color{Black}{\Rightarrow x = y^2 + 2y +4 }\) So, it will be f(x) right? So it's just \(x^2 + 2x + 4\)

OpenStudy (anonymous):

my question says x cubed under a square root.

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

Oh no

Parth (parthkohli):

Can you do it still? I've given a big hint

OpenStudy (anonymous):

i dont know what to do after y+1=sqrt of x^3

Parth (parthkohli):

I'm sleepy *yawns*

OpenStudy (anonymous):

why are you on a math study group then?

Parth (parthkohli):

I'm here to help you, not to give the answers.

OpenStudy (anonymous):

not doing a very good job helping

OpenStudy (anonymous):

Is the following true:\[f(x)=\sqrt{x^3}-1 \]